# Polynomial bounds for surfaces in cusped 3-manifolds

### v2

@@ 2.2. Chunk decomposition
Classical alternating knot and link complements have a well-known decomposition into topological polyhedra. This was suggested by W. Thurston and described by Menasco [30]; see also Lackenby [25] and Purcell [35, Chapters 1 and 11]. A more general decomposition of arborescent knots into angled blocks was defined by Futer and Guéritaud [11], and this was generalized further by Howie and Purcell to weakly generalized alternating links [19]. We briefly review the salient points here; see also [36, Section 3] for further discussion and examples.
@@
    Regions of $\pi(L)$ bounded by interior edges. These are called faces of the chunk. Observe that they are not necessarily simply connected, but they are checkerboard colored.
We further truncate ideal vertices. Because vertices are 4-valent, the resulting truncation turns each ideal vertex into a quad, called a truncation face, bounded by truncation edges (called boundary faces and boundary edges in [19]). Howie and Purcell prove that a weakly generalized alternating link complement admits such a decomposition [19, Propositions 3.1 and 3.3].
@@ 2.3. Normal surfaces
We now consider $(Z,\partial Z)$ to be an essential surface embedded in $(Y-N(L),\partial N(L))$, where the boundary of the surface $Z$ is possibly empty. We will generally allow $Z$ to be orientable or nonorientable, unless otherwise stated. Recall that a surface is essential if it is incompressible, boundary incompressible, and is not boundary parallel.
@@
Throughout, we use the topological definition of incompressible surfaces: a surface $Z$ that is neither a disk nor a 2-sphere is incompressible in a 3-manifold $M$ if any disk $D$ with interior embedded in $M-N(Z)$, with boundary on $Z$, satisfies $\partial D$ bounds a disk in $Z$. A surface that is not incompressible admits an essential compression disk, namely a disk $D$ with interior embedded in $M-N(Z)$ with $\partial D$ an essential curve on $Z$.
@@
We will count essential surfaces later in the paper in weakly generalized alternating links and their Dehn fillings. To do so, we put them into normal form with respect to the chunk decompostion. Normal form in this setting was introduced in Howie–Purcell [19, Definition 3.7]; it generalizes the definition of standard position for surfaces in classical alternating knots used by Menasco [28] and Menasco–Thistlethwaite [31], as well as normal form for surfaces in Lackenby [25] and Futer–Guéritaud [11]. We will not need the full definition of normal form here, only the consequences of that definition from Howie–Purcell [19] and from Purcell–Tsvietkova [36]. Therefore we refer to these references for definitions and examples, but we review the necessary results here.
First, it was shown in Howie–Purcell [19, Theorem 3.8] that any essential surface in a 3-manifold with a chunk decomposition can be put into normal form with respect to the chunk decomposition. When $Z$ is isotoped into normal form, it is cut into components $Z_{i}$, which are connected subsurfaces of $Z$ in normal form within a chunk. These may be closed or with boundary, and might have multiple boundary components. We will write $Z=\bigcup_{i}Z_{i}$.
Both for classical alternating links in $S^{3}$ [28, Theorem 2] and for weakly generalised alternating links under certain conditions [19, Lemma 4.9], a closed surface $Z^{\prime}$ admits meridianal compressions. After meridianal compressions the resulting surface $Z$ can be arranged to meet the diagram in meridians.
It was shown by Purcell–Tsvietkova [36, Theorem 6.2] that an isotopy into normal form can be done such that boundary curves $\partial Z$ that are meridians remain in meridianal form after the isotopy, meaning that $\partial Z$ cuts off exactly two corners of quads corresponding to truncation faces, one on $\Pi^{+}$ and one on $\Pi^{-}$. See Figure 1, which is from [36]. Moreover, such an isotopy does not increase weight, where the weight is the pair $(s(Z),t(Z))$, ordered lexicographically, where $s(Z)$ counts the number of intersections of $Z$ with interior edges, and $t(Z)$ counts the number of intersections with truncation edges.
@@
    out of all ways to isotope the surface into normal, meridianal form, we choose one with least weight.
This allows a labeling of curves of $Z_{i}$ that lie on chunk boundary, described in [36, Section 5], which we now review.
For normal components $Z_{j}$ of the surface $Z$, each boundary component of $\partial Z_{j}$ runs over truncation edges and interior edges of the respective chunks. Each intersection of $\partial Z_{j}$ with an interior edge is labeled with an $S$, for “saddle”, corresponding to an analogous labeling by Menasco [28]; see also [15, 16]. If a component of $\partial Z$ is in meridianal form, then there will be some $Z_{i}$ and $Z_{j}$ such that $\partial Z_{i}$ and $\partial Z_{j}$ each cut off exactly one corner of a truncation quad, in meridianal form. We label these intersections with truncation faces by $P$. Finally, more generally some $\partial Z_{k}$ may run through truncation faces that may or may not be in meridianal form. In this case, it runs through one truncation edge on the way in, and one truncation edge on the way out; label each intersection with a truncation edge with a $B$. The labeling of components of $\partial Z_{i}$ by letters $S$, $B$, or $P$ associates a cyclic word to $\partial Z_{i}$.
For classical alternating links, each $Z_{i}$ can be taken to be a disk lying in one of the topological 3-balls either above or below the link. The boundary curves $\partial Z_{i}$ are determined by the associated words in $S$, $P$, and $B$ and by the position of each letter on the link diagram. Thus bounding the number of possible boundary words gives a bound on the number of surfaces; this is used in the work of Hass, Thompson, and Tsvietkova [15, 16]. For weakly generalized alternating links, the subsurfaces $Z_{i}$ might have positive genus and multiple boundary components. Moreover, the subsurfaces are not in topological 3-balls anymore, but in chunks which might have complicated topology themselves. Thus we need more tools to restrict and control possible subsurfaces that arise. One such tool is the combinatorial area.
@@ 2.4. Combinatorial area
Write $Z=\bigcup_{j=1}^{m}Z_{j}$ where each $Z_{j}$ is a connected normal surface embedded in a chunk. Each curve $\partial Z_{j}$ will meet some number of interior edges, each labeled by $S$; denote the number by (#$(\#S)$. It will also meet truncation edges, with each either labeled by $B$ or with a pair in meridianal form labeled by a single $P$. Denote the number of instances of $B$ by (#$(\#B)$ and the number of instances of $P$ by (#$(\#P)$. On a single component of $N(L)$, all intersections of $\partial Z_{j}$ with corresponding truncation edges will either all be labeled $B$ or all labeled $P$.
@@
This construction satisfies a Gauss–Bonnet formula [19, Proposition 3.12]:
@@
Furthermore, in [36, Lemma 8.5], it was shown that combinatorial area satisfies the following lemma.
@@
The only possible disks with zero area are enumerated in [36, Lemma 8.6]; they have boundaries labeled one of $PP$, $PSS$, $SSSS$, $BBBB$, $BBSS$, or in the case of links, $PBB$. We will always relabel the instance of $P$ in the final case with two instances of $B$; this allows us to identify such disks as $BBBB$ disks and use the tools in that setting to control such disks.
The zero area disks are considered in [36]. By Theorem 9.1 in that paper, $SSSS$ disks only appear as an essential compression disk for $\Pi$, with boundary meeting the diagram $\pi(L)$ in exactly four interior edges. They will not occur if the representativity satisfies $r(\pi(L),\Pi)>4$. By [36, Theorem 10.1], there are no $PP$ disks (nor $PS$ nor $SS$ disks). By [36, Theorem 10.2], there are no $PSS$ disks.
If $Z_{i}$ comes from a spanning surface or more general surface with boundary and meets truncation edges, then there will be letters $B$. We cannot rule these out, but in [36, Section 11] it is shown that if such disks arise, they form larger disks together. In particular, the following appears in that paper.
@@ 3. Meridianal compressions
Let $Z$ be a surface properly embedded in $Y-N(L)$. We say $Z$ is meridianally compressible if there is an essential meridianal annulus $A$ with one component $\partial_{1}A$ of $\partial A$ on $Z$, and the other component $\partial_{2}A$ of $\partial A$ forming a meridian on $\partial N(L)$. The requirement that $A$ be essential means that there is no embedded annulus $A^{\prime}$ in $Z$ such that a component of $\partial A^{\prime}$ agrees with $\partial_{1}A$, the other component of $\partial A^{\prime}$ lies on $\partial Z$, forming a meridianal boundary component on $N(L)$, and $A$ is ambient isotopic to $A^{\prime}$ in $Y-N(L)$ rel $\partial_{1}A$. If there is no essential meridianal annulus for $Z$, we say it is meridianally incompressible. For a meridianally compressible surface, performing surgery along $A$ yields a new surface with boundary forming a meridian of $N(L)$; this is called a meridianal compression of $Z$.
As in [15], we will count surfaces after first performing a maximal number of meridianal compressions. However, unlike [15], we now allow surfaces with additional boundary components and nonorientable surfaces. The next lemmas give bounds on the total number of meridianal compressions we must make. These are in terms of the Euler characteristic of the surface, which combines genus and boundary components. Recall that the genus of the surface is the maximum number of disjoint simple closed curves that can be drawn on the surface without disconnecting it. For orientable surfaces, we will denote the genus by $g_{O}$. For a nonorientable surface, the genus is equal to the number of cross-caps attached to a sphere, and is often called a nonorientable genus. Denote it by $g_{N}$. Then for a closed orientable surface $Z$, $\chi(Z)=2-2g_{O}$, and for a closed nonorientable one, $\chi(Z)=2-g_{N}$. Despite this difference, we will often just use $g$ for genus, meaning both orientable and nonorientable, where this does not affect the calculation. For surfaces with $b$ boundary components, the Euler characteristic is then $\chi(Z)=2-2g_{O}-b$ or $\chi(Z)=2-g_{N}-b$, for the orientable and nonorientable case respectively.
All cases are summarised in the following.
@@
Suppose $Z^{\prime}$ is an essential surface with negative Euler characteristic, with or without boundary, orientable or nonorientable, that is properly embedded in $Y-N(L)$. Let $Z$ be obtained from $Z^{\prime}$ by performing a maximal sequence of meridianal compressions. Then the Euler characteristic of $Z$ is $\chi(Z)=\chi(Z^{\prime})$, and the meridianal compressions have added at most $-4\chi(Z)+2$ boundary components. These are tubed in pairs to obtain $Z^{\prime}$ from $Z$.
Note we could prove a better upper bound of $-2\chi(Z)$ when the surface is closed and orientable, by repeating an argument of [15]. However, since we only need a bound, and not a sharp bound, we state the most general bound to avoid dealing with multiple cases later in the paper.
@@
First, consider one-sided curves. Note that if an annulus $A$ has boundary $\partial_{1}A$ on a one-sided curve, then $A$ must wrap twice around the curve, and gluing $A$ to itself along $\partial_{1}A$ forms a Möbius band with meridianal boundary. However, [19, Theorem 4.6] implies that there is no embedded Möbius band with meridianal boundary in a weakly generalised alternating link. Thus we may rule out one-sided curves for meridianal compressions.
@@
If $Z^{\prime}$ is a closed orientable surface with negative Euler characteristic, then it contains at most $3\displaystyle 3g-$ 3 disjoint nonparallel curves; these cut $Z^{\prime}$ into pairs of pants. More generally, if $Z^{\prime}$ is orientable with genus $g$ and $b$ boundary components, then it contains $3g-3+b$ disjoint nonparallel curves that are not parallel to boundary components. The total number of boundary components after removing the interior of a regular neighborhood of each of these curves is $2\displaystyle 6g-6+2b\leq-3(2-2g-b)=-3\chi(Z^{\prime})<-4\chi(Z^{\prime})+$ 2. This is an upper bound for the number of boundary components produced by meridianal compression in this case.
If $Z^{\prime}$ is a nonorientable surface with $b$ boundary components, then it can be cut along $g$ disjoint nonparallel two-sided curves to cut it into planar surfaces and Möbius bands. An additional $g+b-3$ curves cut the planar surfaces into pairs of pants. This gives $2g-3+b$ disjoint nonparallel two-sided curves. The total number of boundary curves after removing regular neighborhoods of these curves is at most $2\displaystyle 4g-6+2b<4g+4b-8+2=-4\chi(Z)+$ 2. ∎
@@
[36, Proposition 7.1 (1)] There is no meridianal compression of $Z$ to a component $L_{i}$ of $\partial N(L)$ for which $\partial Z\cap L_{i}$ is nonempty and non-meridianal.
@@
From now on, assume $Z$ is essential and properly embedded in $Y-N(L)$. The surface $Z$ is not necessarily orientable, and possibly has boundary components, meridianal or nonmeridianal, that are always on $\partial N(L)$.
@@ 4. Counting subsurfaces and their boundary curves
Let $n_{1}$ be the number of the $Z_{i}$ that have Euler characteristic 00 0 or 11 11, and are not disks with boundary $BBBB$ or $BBSS$. Let $n_{2}$ be the number with Euler characteristic at most $-1$. Then the total number of subsurfaces $Z_{i}$ considered in the lemma is $m=n_{1}+n_{2}$. By the Gauss–Bonnet theorem (2.6), by equation (2.5), and by Lemma 2.7 (1),
@@
For the subsurfaces with $\chi(Z_{i})=1$ or 00 0, we need to determine which have $a(Z_{i})=0$. By Lemma 2.7 (3), these can be tori, Klein bottles, annuli or Möbius bands meeting no edges, or disks meeting four edges. Assumption 2.9, that all regions of $\pi(L)$ on $\Pi$ are disks, rules out annuli and Möbius bands meeting no edges: the boundary curve would lie in one of the cellular faces, and hence bound a disk there, which is not allowed for surfaces in normal form. Assumption 2.9 also requires that $Y-N(\Pi)$ admits no essential tori and Klein bottles, which rules out these surfaces. Thus if $a(Z_{i})=0$ then $Z_{i}$ is a normal disk meeting four edges.
If $Z$ has only meridianal boundary components and $a(Z_{i})=0$, then it is a disk with boundary of the form $PSS$, $PP$, or $SSSS$. But these are ruled out by [36, Theorem 10.1], [36, Theorem 10.2], and [36, Theorem 9.1], respectively, using the hypothesis that $r(\pi(L),\Pi)>4$ for the $SSSS$ case. If $Z$ has a non-meridianal boundary component, then we consider all intersections with truncation edges as labeled $B$. The only remaining normal disks with $a(Z_{i})=0$ are $BBBB$ or $BBSS$ disks, ruled out by hypothesis.
@@
$\sum_{\chi(Z_{i})=0,1}a(Z_{i})\geq n_{1}\,\frac{\pi}{2},$
and hence
[…]
@@
$-2\pi\chi(Z)\>=\>a(Z)\>=\>\sum_{i=1}^{m}a(Z_{i}).$
Moreover, each time a boundary component of $Z$ meets an interior or truncation edge of the chunk decomposition, there is a contribution of $\pi/2$ to the area. More precisely, if $w_{i}$ is the total number of intersections of all the boundary components of $Z_{i}$ with chunk edges, then
@@
In case (1), when the surface has meridianal boundary, all words are in letters $P$ and $S$, so $w=\sum w_{i}$. In case (2), when $Z_{i}$ is a $BBBB$ or $BBSS$ disk, we know its combinatorial area is 00 0. Moreover, we know that $2\pi\chi(Z_{i})\leq 0$ unless $Z_{i}$ is a disk. Thus from equation (4.4), in both cases we have
$a(Z)\>\geq\>w\frac{\pi}{2}-2\pi\cdot\left|\{Z_{i}:Z_{i}\mbox{ isa disk, but not a $\displaystyle BBBB$ or $\displaystyle BBSS$ disk}\}\right|.$

### v3

@@ 2.2. Chunk decomposition
Classical alternating knot and link complements have a well-known decomposition into topological polyhedra. This was suggested by W. Thurston and described by Menasco [43]; see also Lackenby [38] and Purcell [48, Chapters 1 and 11]. A more general decomposition of arborescent knots into angled blocks was defined by Futer and Guéritaud [20], and this was generalized further by Howie and Purcell to weakly generalized alternating links [27]. We briefly review the salient points here; see also [49, Section 3] for further discussion and examples.
@@
    Regions of $\pi(L)$ bounded by interior edges. These are called faces of the chunk. They are not necessarily simply connected, but they are checkerboard colored.
We further truncate ideal vertices. Because vertices are 4-valent, the resulting truncation turns each ideal vertex into a quad, called a truncation face, bounded by truncation edges (called boundary faces and boundary edges in [27]). A weakly generalized alternating link complement admits such a decomposition [27, Propositions 3.1 and 3.3].
@@ 2.3. Normal surfaces
We now consider $(Z,\partial Z)$ to be an essential surface embedded in $(Y-N(L),\partial N(L))$, where the boundary of the surface $Z$ is possibly empty. We generally allow $Z$ to be orientable or nonorientable, unless otherwise stated. Recall a surface is essential if it is incompressible, boundary incompressible, and is not boundary parallel.
@@
Throughout, we use the topological definition of incompressible surfaces: a surface $Z$ that is neither a disk nor a 2-sphere is incompressible in a 3-manifold $M$ if any disk $D$ with interior embedded in $M-N(Z)$, with boundary on $Z$, satisfies $\partial D$ bounds a disk in $Z$. A surface that is not incompressible admits an essential compression disk, i.e. a disk $D$ with interior embedded in $M-N(Z)$ with $\partial D$ an essential curve on $Z$.
@@
We will count essential surfaces later in the paper in weakly generalized alternating links and their Dehn fillings. To do so, we put them into normal form with respect to the chunk decompostion. Normal form in this setting was introduced in Howie–Purcell [27, Definition 3.7]; it generalizes the definition of standard position for surfaces in classical alternating knots used by Menasco [41] and Menasco–Thistlethwaite [44], as well as normal form for surfaces in Lackenby [38] and Futer–Guéritaud [20]. We will not need the full definition of normal form here, only the consequences of that definition from Howie–Purcell [27] and from Purcell–Tsvietkova [49]. Therefore we refer to these references for definitions and examples, but we review the necessary results here.
In  [27, Theorem 3.8], it is shown that any essential surface in a 3-manifold with a chunk decomposition can be put into normal form with respect to the chunk decomposition. When $Z$ is isotoped into normal form, it is cut into components $Z_{i}$, which are connected subsurfaces of $Z$ in normal form within a chunk. These may be closed or with boundary, and might have multiple boundary components. We will write $Z=\bigcup_{i}Z_{i}$.
Both for classical alternating links in $S^{3}$ [41, Theorem 2] and for weakly generalized alternating links under certain conditions [27, Lemma 4.9], a closed surface $Z^{\prime}$ admits meridianal compressions. After meridianal compressions the resulting surface $Z$ can be arranged to meet the diagram in meridians.
In [49, Theorem 6.2], it is shown that an isotopy into normal form can be done such that boundary curves $\partial Z$ that are meridians remain in meridianal form after the isotopy, meaning that $\partial Z$ cuts off exactly two corners of quads corresponding to truncation faces, one on $\Pi^{+}$ and one on $\Pi^{-}$. See Figure 1, which is from [49]. Moreover, such an isotopy does not increase weight, where the weight is the pair $(s(Z),t(Z))$, ordered lexicographically, where $s(Z)$ counts the number of intersections of $Z$ with interior edges, and $t(Z)$ counts the number of intersections with truncation edges.
$\Pi^{+}$
$\Pi^{-}$
@@
    of all ways to isotope the surface into normal, meridianal form, it has least weight.
We obtain a labeling of curves of $Z_{i}$ that lie on the chunk boundary, described in [49, Section 5], which we now review. For a normal component $Z_{j}$ of the surface $Z$, each boundary component of $\partial Z_{j}$ runs over truncation edges and interior edges of a chunk. Each intersection of $\partial Z_{j}$ with an interior edge is labeled with an $S$ for “saddle”, corresponding to an analogous labeling by Menasco [41]; see also [24, 25]. If a component of $\partial Z$ is in meridianal form, there will be some $Z_{i}$ and $Z_{j}$ such that $\partial Z_{i}$ and $\partial Z_{j}$ each cut off exactly one corner of a truncation quad in meridianal form. Label these intersections with faces by $P$. More generally some $\partial Z_{k}$ may run through truncation faces that may or may not be in meridianal form. Then $\partial Z_{k}$ intersects one truncation edge on the way in, and one truncation edge on the way out; label each intersection with a $B$. The labeling of components of $\partial Z_{i}$ by letters $S$, $B$, or $P$ associates a cyclic word to $\partial Z_{i}$.
For classical alternating links, each $Z_{i}$ can be taken to be a disk lying in one of the topological 3-balls either above or below the link. The boundary curves $\partial Z_{i}$ are determined by the associated words in $S$, $P$, and $B$ and by the position of each letter on the link diagram. Thus bounding the number of possible boundary words gives a bound on the number of surfaces; this is used in the work of Hass, Thompson, and Tsvietkova [24, 25]. For weakly generalized alternating links, the subsurfaces $Z_{i}$ might have positive genus and multiple boundary components. Moreover, the subsurfaces are not in topological 3-balls anymore, but in chunks which might have complicated topology themselves. Hence we need more tools to restrict and control possible subsurfaces that arise. One such tool is the combinatorial area.
@@ 2.4. Combinatorial area
Write $Z=\bigcup_{j=1}^{m}Z_{j}$ where each $Z_{j}$ is a connected normal surface embedded in a chunk. Each curve $\partial Z_{j}$ will meet some number of interior edges, each labeled by $S$; denote the number by $(\#S)$. It will also meet truncation edges, with each either labeled by $B$ or with a pair in meridianal form labeled by a single $P$. Denote the number of instances of $B$ by $(\#B)$ and the number of instances of $P$ by $(\#P)$. On a single component of $N(L)$, all intersections of $\partial Z_{j}$ with corresponding truncation edges will either all be labeled $B$ or all labeled $P$.
@@
This construction satisfies a Gauss–Bonnet formula [27, Proposition 3.12]:
@@
Furthermore, combinatorial area satisfies the following lemma.
@@
The only possible disks with zero area are enumerated in [49, Lemma 8.6]; they have boundaries labeled one of $PP$, $PSS$, $SSSS$, $BBBB$, $BBSS$, or in the case of links, $PBB$. The instance of $P$ in the final case will be relabeled with two instances of $B$; this allows us to identify such disks as $BBBB$ disks and use tools in that setting.
The zero area disks are considered in [49]. By Theorem 9.1 in that paper, $SSSS$ disks only appear as an essential compression disk for $\Pi$, with boundary meeting the diagram $\pi(L)$ in exactly four interior edges. They will not occur if the representativity satisfies $r(\pi(L),\Pi)>4$. By [49, Theorem 10.1], there are no $PP$ disks (nor $PS$ nor $SS$ disks). By [49, Theorem 10.2], there are no $PSS$ disks.
If $Z_{i}$ comes from a spanning surface or more general surface with boundary and meets truncation edges, then there will be letters $B$. We cannot rule these out, but in [49, Section 11] it is shown that if such disks arise, they form larger disks together. In particular, the following appears in that paper.
@@ 3. Meridianal compressions
Let $Z$ be a surface properly embedded in $Y-N(L)$. We say $Z$ is meridianally compressible if there is an essential meridianal annulus $A$ with one component $\partial_{1}A$ of $\partial A$ on $Z$, and the other component $\partial_{2}A$ of $\partial A$ forming a meridian on $\partial N(L)$. The requirement that $A$ be essential means it is not ambient isotopic rel $\partial_{1}A$ to an embedded annulus $A^{\prime}$ in $Z$ with $\partial A^{\prime}$ the disjoint union of $\partial_{1}A$ and a meridian of $N(L)$. If there is no essential meridianal annulus for $Z$, we say it is meridianally incompressible. For a meridianally compressible surface, performing surgery along $A$ yields a new surface with boundary forming a meridian of $N(L)$; this is called a meridianal compression of $Z$.
As in [24], we will count surfaces after first performing a maximal number of meridianal compressions. However, unlike [24], we allow surfaces with additional boundary components and nonorientable surfaces. The next lemmas give bounds on the total number of meridianal compressions we must make, in terms of the Euler characteristic $\chi$ of the surface. Recall that the genus of the surface is the maximum number of disjoint simple closed curves that can be drawn on the surface without disconnecting it. For orientable surfaces, we denote the genus by $g_{O}$. For a nonorientable surface, the genus is equal to the number of cross-caps attached to a sphere; it is often called a nonorientable genus. Denote it by $g_{N}$. Then for a closed orientable surface $Z$, $\chi(Z)=2-2g_{O}$, and for a closed nonorientable one, $\chi(Z)=2-g_{N}$. For surfaces with $b$ boundary components, the Euler characteristic is $\chi(Z)=2-2g_{O}-b$ or $\chi(Z)=2-g_{N}-b$, for the orientable and nonorientable case respectively. We often denote genus by $g$, for both orientable and nonorientable surfaces, where this does not affect the calculation.
@@
Suppose $Z^{\prime}$ is an essential surface with negative Euler characteristic, with or without boundary, orientable or nonorientable, that is properly embedded in $Y-N(L)$. Let $Z$ be obtained from $Z^{\prime}$ by performing a maximal sequence of meridianal compressions. Then the Euler characteristic of $Z$ is $\chi(Z)=\chi(Z^{\prime})$, and the meridianal compressions add at most $-4\chi(Z)+2$ boundary components. These are tubed in pairs to obtain $Z^{\prime}$ from $Z$.
Note we could prove a better upper bound of $-2\chi(Z)$ when the surface is closed and orientable, by repeating an argument of [24]. However, since we only need a bound, and not a sharp bound, we state the most general bound to avoid multiple cases later.
@@
First, consider one-sided curves. Note that if an annulus $A$ has boundary $\partial_{1}A$ on a one-sided curve, then $A$ must wrap twice around the curve, and gluing $A$ to itself along $\partial_{1}A$ forms a Möbius band with meridianal boundary. However, [27, Theorem 4.6] implies that there is no embedded Möbius band with meridianal boundary in a weakly generalized alternating link. Thus we may rule out one-sided curves for meridianal compressions.
@@
If $Z^{\prime}$ is a closed orientable surface with negative Euler characteristic, it contains at most $3g-3$ disjoint nonparallel curves; these cut $Z^{\prime}$ into pairs of pants. More generally, if $Z^{\prime}$ is orientable with genus $g$ and $b$ boundary components, it contains $3g-3+b$ disjoint nonparallel curves that are not parallel to boundary components. The number of boundary components after removing a regular neighborhood of each of these curves is $6g-6+2b\leq-3(2-2g-b)=-3\chi(Z^{\prime})<-4\chi(Z^{\prime})+2$. This is an upper bound for the number of boundary components produced by meridianal compression in this case.
If $Z^{\prime}$ is a nonorientable surface with $b$ boundary components, it can be cut into planar surfaces and Möbius bands along $g$ disjoint nonparallel two-sided curves. An additional $g+b-3$ curves cut the planar surfaces into pairs of pants, giving $2g-3+b$ disjoint nonparallel two-sided curves. The total number of boundary curves after removing regular neighborhoods of these curves is at most $4g-6+2b<4g+4b-8+2=-4\chi(Z)+2$. ∎
@@
[49, Proposition 7.1 (1)] There is no meridianal compression of $Z$ to a component $L_{i}$ of $\partial N(L)$ for which $\partial Z\cap L_{i}$ is nonempty and non-meridianal.
@@
From now on, assume $Z$ is essential, properly embedded in $Y-N(L)$. The surface $Z$ is not necessarily orientable, and possibly has boundary components, meridianal or nonmeridianal, that are always on $\partial N(L)$.
@@ 4. Counting subsurfaces and their boundary curves
Let $n_{1}$ be the number of the $Z_{i}$ that have Euler characteristic $0$ or $1$, and are not disks with boundary $BBBB$ or $BBSS$. Let $n_{2}$ be the number with Euler characteristic at most $-1$. Then the total number of subsurfaces $Z_{i}$ considered in the lemma is $m=n_{1}+n_{2}$. By the Gauss–Bonnet theorem (2.6), by equation (2.5), and by Lemma 2.7 (1),
@@
For the subsurfaces with $\chi(Z_{i})=1$ or $0$, we need to determine which have $a(Z_{i})=0$. By Lemma 2.7 (3), these can be tori, Klein bottles, annuli or Möbius bands meeting no edges, or disks meeting four edges. Assumption 2.9, that all regions of $\pi(L)$ on $\Pi$ are disks, rules out annuli and Möbius bands meeting no edges: the boundary curve would lie in one of the cellular faces, and hence bound a disk there, which is not allowed for surfaces in normal form. Assumption 2.9 also requires that $Y-N(\Pi)$ admits no essential tori and Klein bottles, which rules out these surfaces. Thus if $a(Z_{i})=0$ then $Z_{i}$ is a normal disk meeting four edges.
If $Z$ has only meridianal boundary components and $a(Z_{i})=0$, then it is a disk with boundary of the form $PSS$, $PP$, or $SSSS$. But these are ruled out by [49, Theorem 10.1], [49, Theorem 10.2], and [49, Theorem 9.1], respectively, using the hypothesis that $r(\pi(L),\Pi)>4$ for the $SSSS$ case. If $Z$ has a non-meridianal boundary component, then we consider all intersections with truncation edges as labeled $B$. The only remaining normal disks with $a(Z_{i})=0$ are $BBBB$ or $BBSS$ disks, ruled out by hypothesis.
@@
$\sum_{\chi(Z_{i})=0,1}a(Z_{i})\geq n_{1}\,\frac{\pi}{2},\quad\mbox{and hence}$
$2\pi\chi(Z)\geq n_{1}\,\frac{\pi}{2}+n_{2}\,2\pi\geq\frac{\pi}{2}\,m.$
@@
$2\pi\chi(Z)\>=\>a(Z)\>=\>\sum_{i=1}^{m}a(Z_{i}).$
Each time a boundary component of $Z$ meets an interior or truncation edge of the chunk decomposition, there is a contribution of $\pi/2$ to the area. More precisely, if $w_{i}$ is the total number of intersections of all the boundary components of $Z_{i}$ with chunk edges, then
@@
In case (1), when the surface has meridianal boundary, all words are in letters $P$ and $S$, so $w=\sum w_{i}$. In case (2), when $Z_{i}$ is a $BBBB$ or $BBSS$ disk, we know its combinatorial area is $0$. Moreover, we know that $2\pi\chi(Z_{i})\leq 0$ unless $Z_{i}$ is a disk. Thus from equation (4.4), in both cases we have
[…]